magnetic field is 37.2..at the center of the circular coil of radius 42cm..that has 100 turns  is

Grade:12

2 Answers

SHANMUKESHWAR
8 years ago
radius=42*10^-2
B=37.2
turns=100
B=4*3.14*I*100/2*42*10^-2
B=I*0.149*10^-3
I=37.2/0.149*10^-3
I=2.49*10^5A
Prabhakar ch
8 years ago
Dear Sai kumar
                         Radius, R=42*10^-2
                         Magnetic field,  B=37.2
                         turns=100
                         B=4*3.14*I*100/2*42*10^-2
                         B=I*0.149*10^-3
                         I=37.2/0.149*10^-3
                         I=2.49*10^5A